OpenStudy (anonymous):

I've been stuck on this problem for the an hour and a half... please help. Find the solution to the equation log16^32 = x + 5. Using complete sentences, explain the procedure used to solve this equation.

OpenStudy (anonymous):

log16/log32 = x + 5

OpenStudy (anonymous):

How'd you get that? :O

OpenStudy (anonymous):

Formula logb^U = logb/logU from Pre-Calc

OpenStudy (anonymous):

Well this is algebra 2 so they haven't taught us that :/ They only gave us this formula: y = b^x x = logb y

OpenStudy (unklerhaukus):

\[\log_{16}{32} = x + 5\]\[\frac{\log_232}{\log_216}=x+5\]\[\frac{\log_22^5}{\log_22^4}=x+5\]\[\frac{5\log_22}{4\log_22}=x+5\]\[\frac{5}{4}=x+5\]\[\cdots=x\]

OpenStudy (anonymous):

then it should be 16^(x+5) = 32 because ( log16^32 = x+5) use the x=logb y => y=b^x 2^4(x+5)= 2^5 4(x+5) = 5 ( omit 2 because they have the same base) 4x + 20 = 5 4x = -15 x = -15/4

OpenStudy (anonymous):

Ohh, sorry it's hard learning off of a virtual teacher. Thank you, i think i get it now